OpenStudy (anonymous):

Solve the equation: 1/2(4x-6)=-3x+2 Please show steps, not just answer thank you.

OpenStudy (diyadiya):

1/2(4x-6)=-3x+2 Multiply both sides by 2 what do you get ?

OpenStudy (diyadiya):

\[ \frac{1}{2}(4x-6) \times 2 = (-3x+2) \times 2\]

OpenStudy (diyadiya):

\[ \frac{1}{\cancel2}(4x-6)\times {\cancel2}= (-3x+2) \times 2\]

OpenStudy (anonymous):

I'm sorry can you put this in numbers without all the symbols, please and thank you so much for helping me.

OpenStudy (diyadiya):

Without using the equation editor?

OpenStudy (anonymous):

Ok that's much better, thanks

OpenStudy (diyadiya):

Can you try it out ? 1/2(4x-6)=-3x+2 first multiply both sides by 2

OpenStudy (diyadiya):

Now tell me what you got after multiplying ? Its ok if you're doing it wrong

OpenStudy (diyadiya):

I can help you

OpenStudy (anonymous):

how do I do it with that fraction at the beginning?

OpenStudy (diyadiya):

 Now as you can see the 2 will cancel out on the left side